Fermi National Laboratory
 

Operation Summary for December 21, 2009 - January 4, 2010

 

The Integrated Luminosity for the period from 12/21/09 to 1/4/10 was 122.1 pb-1.
NuMI reported receiving 17.37E18 protons on target during this same period.

Monday Dec. 21

The day shift began with Operations monitoring Tevatron (TeV) store 7447 (156.01E30) and the stash (123.8E10) and the Antiproton Source stack (22.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, NuMI users took beam smoothly except for a short half-hour Absorber Hall access in the afternoon. MTest users took beam all day.

During the evening shift, Operations terminated store 7447 at 8 PM and established store 7448 at 10 PM with an average initial luminosity (AIL) of 269E30. Operations disabled the MTest beam at 6 PM.

Tuesday Dec. 22

The midnight shift began with Operations monitoring store 7448 (165.16E30) and the stash (105.8E10), with Pbar stacking (23.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations resumed sending beam to MTest at 4:30 AM.

During the day shift, Operations terminated store 7448 at 1 PM and established store 7449 at 2:30 PM with an average initial luminosity (AIL) of 297.8E30. NuMI personnel made a short access into their Absorber Hall.

During the evening shift, Operations disabled the MTest beam at 8:15 PM.

Wednesday Dec. 23

The midnight shift began with Operations monitoring store 7449 (100.0E30) and the stash (270.0E10), with Pbar stacking (22.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7449 at 6 AM.

During the day shift, Operations established store 7450 with an AIL of 291.4E30. Late in the morning a H- Source dome access was made and Pbar target studies were performed in the afternoon.

During the evening shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available. Pbar transfers to the Recycler were completed when appropriate.

Thursday Dec. 24

The midnight shift began with Operations monitoring store 7450 (65.09E30) and the stash (358.9E10), with Pbar stacking (36.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7450 was terminated at 3 AM; shot setup followed and concluded with store 7451 with an AIL of 308.6E30.

During the day shift, time was spent from 9:30 until 10 AM repairing Booster and Main Injector RF stations.

During the evening shift, store 7451 was terminated at 10:30 PM and was followed by shot setup which yielded store 7452 with its AIL of 303E30 by 11:15 PM.

Friday Dec. 25

The midnight shift began with Operations monitoring store 7452 (240.04E30) and the stash (73.4E10), with Pbar stacking (20.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7452 was terminated as scheduled at 3:30 AM; shot setup followed and continued into the next shift. A short amount of downtime was accrued due to BRF5 and L:QPS104.

During the day shift, shot setup continued from the previous shift and store 7453 was colliding with an AIL of 323.7E30 by 4:30 PM.

During the evening shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available with Pbar transfers to the Recycler completed when appropriate.

Saturday Dec. 26

The midnight shift began with Operations monitoring store 7453 (120.77E30) and the stash (210.8E10), with Pbar stacking (24.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available, and with Pbar transfers to the Recycler completed when appropriate.

During the day shift, store 7453 collided until the scheduled termination time of 10:15 AM; shot setup followed and store 7454 was colliding with an AIL of 288E30 by 11:54 AM.

During the evening shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available, and with Pbar transfers to the Recycler completed when appropriate.

Sunday Dec. 27

The midnight shift began with Operations monitoring store 7454 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7454 collided until the scheduled termination time of 5:10 AM; shot set-up followed and store 7455 was colliding with an AIL of 286E30 by 6:50 AM. An Autotune setting problem with ECool correctors inhibited one final antiproton transfer to the Recycler and delayed shot setup by a small amount in order to cool the stash. Recycler personnel were contacted and the correctors were recovered.

During the day shift, Pbars were stacked and transferred to the Recycler as appropriate. Beam was delivered to NuMI and MiniBooNE when it was available.

During the evening shift, store 7455 collided until the scheduled termination time of 9:45 PM; shot setup followed and store 7456 was colliding with an AIL of 278E30 by 11:45 PM.

Monday Dec. 28

The midnight shift began with Operations monitoring store 7457 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, a networking issue with a Pbar cooling system PLC was resolved.

During the day shift, MI proton coalescing was checked and Pbar personnel conducted target beam studies.

During the evening shift, store 7456 was terminated as scheduled at 4:15 PM; shot setup followed and store 7457 was colliding with an AIL of 300E30 by 6:15 PM. The Pbar lithium lens tripped multiple times; target station personnel were contacted and instructions for future trips were developed.

Tuesday Dec. 29

The midnight shift began with Operations monitoring store 7457 (129.43E30) and the stash (185.1E10), with Pbar stacking (21.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations monitored and increased the Booster LLRF temperature appropriately. The Pbar lens tripped again and the offending interlock hardware was jumpered per Target personnel instructions.

During the day shift, store 7457 collided until the scheduled termination time of 11 AM; shot setup followed and store 7458 was colliding with an AIL of 314E30 by 12:45 PM. Booster personnel made an access to pull the BRF19 cavity out of the enclosure. Other small repairs were made during the downtime.

During the evening shift, instructions were left with the next crew to address issues with T:HELIX between stores.

Wednesday Dec. 30

The midnight shift began with Operations monitoring store 7458 (85.59E30) and the stash (259.4E10), with Pbar stacking (25.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7458 collided until the scheduled termination time of 7:00 AM; shot set-up followed and continued into the next shift.

During the day shift, Operations established store 7459 with an AIL of 316E30 by 8:45 AM. Pbar target station personnel conducted 2 hours of beam studies.

During the evening shift, store 7459 was terminated as scheduled at 9 PM; collider hall accesses followed. TeV turn around and orbit checks continued into the next shift. Operators monitored the I- Source.

Thursday Dec. 31

The midnight shift began with Operations monitoring the stash (376.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, shot setup began by 1:15 AM; would be store 7463 quenched at 2:20 AM following the low beta squeeze, thus ending shot setup. The quench investigation and recovery followed. No obvious hardware problems were uncovered. The machine was checked out and turned back on. TeV beam studies began at 6 AM and continued into the next shift.

During the day shift, TeV ramp/squeeze studies lasted until 10:45 AM; shot setup followed and store 7467 was colliding with an AIL of 219.75E30 by 12:30 PM.

During the evening shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available with Pbar transfers to the Recycler completed when appropriate.

Friday Jan. 1

The midnight shift began with Operations monitoring store 7467 (63.32E30) and the stash (308.6E10), with Pbar stacking (22.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7467 was terminated at 4:51 AM; shot setup followed and store 7468 was colliding with an AIL of 303.3E30 at 6:10 AM.

During the day shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available with Pbar transfers to the Recycler completed when appropriate.

During the evening shift, store 7468 was terminated at 9:59 PM; shot setup followed and store 7469 was colliding with an AIL of 306.7E30 at 11:54 PM.

Saturday Jan. 2

The midnight shift began with Operations monitoring store 7469 (289.45E30) and the stash (63.8E10), with Pbar stacking (13.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available with Pbar transfers to the Recycler completed when appropriate.

During the day shift, MiniBooNE did not take beam from 8:38 to 11:55 AM. Store 7469 was terminated at 2:59 PM; shot setup followed and store 7470 was colliding with an AIL 312.81E30 at 4:58 PM.

During the evening shift, beam was delivered to the Pbar Source, NuMI, and MiniBooNE when available with Pbar transfers to the Recycler completed when appropriate.

Sunday Jan. 3

The midnight shift began with Operations monitoring store 7470 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TeV quenched at 2:49 AM due to a kicker prefire on T:AAK5, killing store 7470.

During the day shift, recovery of the cryo system followed and continued until 12:37 PM. TeV Turn-on/Check-out and shot setup followed.

During the evening shift, the TeV aborted during shot setup at 6:59 PM due to D0 SVX beam positions and because of a T:QC42 failure. The regulator for T:QC42 was replaced by 10:26 PM and a TeV checkout was performed.

Monday Jan. 4

The midnight shift began with Operations monitoring the stash (185.7E10), with Pbar stacking (25.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into shot setup and store 7477 was colliding with an AIL of 145.6E30 at 1:35 AM.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/04/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory